Can you help me solve the solution of this set of equations x^2+xy+y^2=12 and x+y=2
Thank you

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Can you help me solve the solution of this set of equations x^2+xy+y^2=12 and x+y=2
----------
y = 2-x
Sub for y in the 1st eqn, solve for x.
etc

Answer by stanbon(75887)

Can you help me solve the solution of this set of equations
x^2+xy+y^2=12 and x+y=2
---------
y = 2-x
---
Substitute for "y" and solve for "x":
x^2 + x(2-x) + (2-x)^2 = 12
----
x^2 + 2x - x^2 + 4-4x+x^2 = 12
-----
x^2 -2x + 4 = 12
-----
x^2 -2x - 8 = 0
-------
Factor::
(x-4)(x+2) = 0
----
x = 4 or x = -2
----
If x = 4, y = 2-x = -2
If x = -2 y = 2-x = 4
----------------------------
Cheers,
Stan H.
---------------------